Clermont May 28

Clermont’s bees continue to store honey and produce brood. The BroodMinder™ equipped colony is reporting a total colony weight of 212 pounds. This colony started the season at about 105 pounds. Some of that weight gain is the two ten-frame supers of foundation added as the nectar flow got underway. The foundation is mostly drawn out now and being capped as the flow continues. This colony has brood in egg and larva stages. There was no sign of queen activity at our last visit on May 14th, so a presumption is that the colony swarmed and now has a new active queen. There is reasonable bee coverage in the colony.  

Most of the other colonies have brood, productive queens, and/or queen cells. Because it is not known which colonies may have swarmed, the queen cells were not intentionally disturbed. All the colonies are four supers minimum, and most are five supers. Colonies had substantial orange pollen being brought in by returning field bees. There is no way to observe how much nectar is flowing but inferred evidence of a continuing and substantial nectar flow from the capped and nectar-filled frames is apparent. There are abundant blooming plants visible around the apiary.

Five frame plywood nucs from April were still at the apiary. These nucs are from Clermont swarms, swarms from other yards, or purchased queens made into nucs on April 11th.  All but one of these colonies were out of space with all frames filled, and very limited brood space. These colonies were moved into two supers of five-frame nuc equipment for a total of ten frames. Five frames of foundation and drawn comb frames, split between both supers were utilized to alleviate the space issue. These nucs will be overwintered in a thirteen-frame, three super configuration if not used for enhancing other colonies. One weak nuc was left in a single super configuration. It has evidence of brood but not in a good pattern.

Several colonies exhibited evidence of increasing mite infestation with spotty brood and poor larva conditions. One colony was tested with an alcohol wash and found to have about a 6% mite level. The sample count was eighteen mites from the three hundred bee sample. All the colonies were treated with Formic Pro earlier this season which suggests some colonies were not properly controlled. There was an in-depth discussion on what to do ranging from do nothing while honey supers are on to remove the honey supers and treat with products labeled for warmer temperature. After some consideration, a review of the next ten-day forecast and the availability of the treatment product, the plan is to treat with Formic Pro™ next week on Thursday June 2nd. Monday – Wednesday is forecast to be above the 85⁰F maximum allowed first three-day temperature. June 2nd through June 11th is forecast to be below the upper temperature limit. The maximum ten-day dose will be used on all the standard size colonies that are logged as having possible mite issues. There is the possibility of colony damage from treating with formic acid, but a 6% mite count is considered too high to wait until honey supers are off in late July.